# Jump Starting Bandits with LLM-Generated Prior Knowledge

### v1

@@ 2 Related Work
Recently, researchers have attempted to integrate LLMs into social science studies, specifically utilizing LLMs to simulate human behaviours in social psychology experiments [17, 2] predict their opinions [6, 40]. Additionally, a growing body of research has focused on evaluating LLMs’ performance and biases through survey questions. Prior studies have explored how LLMs respond to multiple-choice questions from public opinion polls, revealing insights into their alignment with human responses and potential biases [39, 19, 18]. Santurkar et al. [39] focus on diverse demographic groups across the U.S., while Durmus et al. [19] focus on non-American demographics, both finding that LLMs’ responses to surveys show discrepancies with the collected answers from representative samples. Furthermore, Dominguez-Olmedo et al. [18] show that models’ responses are sensitive to the ordering and labelling biases, and they do not contain the statistical signals typically found in human populations.
@@ 3 Preliminaries
In a multi-armed bandit problem, an agent (the “bandit”), at each step $t\in[T]$, is presented with a set of $K$ possible arms and must choose an arm $k_{t}$. After which the agent receives a reward $r_{t,k_{t}}$ drawn from a reward distribution $R_{k_{t}}$, which is initially unknown. We focus on the stochastic setting, where each $r_{t,k}$ is sampled independently from $R_{k}$. The goal of the agent is to maximize the total reward $\sum_{t=1}^{T}r_{t,k_{t}}$. To achieve this, the agent needs to balance the trade-off between exploration (i.e., trying different arms to gather more information about their reward distribution) and exploitation (i.e., choosing among the arms that appear to offer the highest reward based on the collected information so far).
Contextual Multi-Arm Bandits.
In contextual multi-arm bandits (CB), where at each time step $t\in[T]$, the agent receives side information $\phi_{t}\in\mathcal{C}$ which is called the context. The reward distribution of pulling arm $k\in[K]$ at time $t$ may be different given the context $\phi_{t}$, that is, $r_{t,k}\sim R_{k}(\cdot\mid\phi_{t})$.
@@
A policy is denoted by $\pi\in\Delta^{K}$ where $\Delta^{K}\subseteq\mathbb{R}^{K}$ is the standard $K$-simplex. A policy $\pi$ chooses arm $k$ given the context $\phi$ with probability of $\pi(k)$. The expected reward of agent at step $t$ under policy $\pi$ given the context $\phi_{t}$ is $r^{\pi}_{t}=\sum_{k=1}^{K}\pi(k)\mathbb{E}\left[r_{t,k}|\phi_{t}\right]$.
As the bandit performs online learning, the policy distribution changes with $t\in[T]$. So the CB algorithm $\mathcal{A}$ has a policy $\pi_{t}\in\Delta^{K}$ and samples $k_{t}\in[K]$ according to $\pi_{t}$, and then the agent observes the corresponding reward $r_{t,k_{t}}$.
@@
The (cumulative) regret after $T$ steps due to the agent choosing policies based on an algorithm $\mathcal{A}$ as $\pi_{1},\pi_{2},\cdots,\pi_{T}$ compared to policies of algorithm $\mathcal{A}^{*}$ as $\pi^{*}_{1},\pi^{*}_{2},\cdots,\pi^{*}_{T}$ is defined is
$\operatorname{reg}_{\mathcal{A},\mathcal{A}^{*}}=\sum_{t=1}^{T}r_{t,k^{*}_{t}}-\sum_{t=1}^{T}r_{t,k_{t}}$
where $k_{t}$ is chosen based on policy $\pi_{t}$ and $k^{*}_{t}$ is chosen based on policy $\pi^{*}_{t}$.
@@
In some applications, not all actions are available at all times. In this sleeping bandit setting [26], the learning algorithm needs to learn and determine which arms yield the best reward when available.
@@ 4.1 Problem Formulation
We consider a set of $n$ users, where the vector of features of each user $i\in[n]$ is denoted by $X_{i}$, which is sampled from a population $\mathcal{X}$. At each time step $t$, user $u_{t}$ is sampled from the set of $n$ users. The user’s context can be calculated using a mapping function $\phi:\mathcal{X}\rightarrow\mathcal{C}$ where $\mathcal{C}$ is the context space. There are $K$ different arms, each representing a potential recommendation for the user. For each user $u_{t}$ and arm $k$, there is a scalar reward $R_{u_{t}}(k)$ following some unknown distribution. For brevity, we write $R_{t}$ henceforth. We note that there is an optimal policy $\pi_{o}$ which chooses the optimal arm $k^{\ast}_{t}=\operatorname*{arg\,max}_{k\in[K]}\mathbb{E}[R_{t}(k)]$ at each time step. We seek to optimize the following (expected, cumulative) regret:
$\pi^{*}=\arg\min_{\pi}\sum_{t=1}^{T}\mathbb{E}_{k_{t}\sim\pi_{o}(X_{t})}\mathbb{E}\left[R_{t}(k_{t})\right]$
$-\sum_{t=1}^{T}\mathbb{E}_{k_{t}\sim\pi(X_{t})}\mathbb{E}\left[R_{t}(k_{t})\right]$
@@ 5 Experiments
We validate the effectiveness of our proposed approach through two sets of experiments. In both sets of experiments, we adapt the LinUCB algorithm [13] to train our bandit. The first set of experiments is based on synthetic data, and the second set uses a real-world dataset derived from a choice-based conjoint analysis [28]. Experimental details are in Appendix C.
@@ 5.1 First Experiment: Personalized Email Campaign for Charity Donations
For this experiment, the ideal participants would be actual donors of our target charity. These users would be characterized by a rich set of attributes such as age, gender, location, occupation, hobbies, and their history of charitable donations. However, obtaining such comprehensive and detailed user data is not only challenging due to the sensitivity and privacy concerns associated with handling personal information but also financially and operationally demanding. Therefore, in this set of experiments, we propose an innovative alternative. Instead of relying on actual user datasets, we utilize a larger model’s responses as an oracle to while pre-training our bandit on the responses of a smaller model. In this experiment, we rely on OpenAI GPT-4o model [1] as a surrogate to the real potential donors, and simulate their preferences.
@@ 7 Limitations
Finally, this work has focused on cases where there are relatively few arms for the bandit to choose from. Further work is necessary to determine how well these results scale to a larger number of arms. Moreover, our reward function estimation focuses on determining the rank order of rewards and not the magnitude of rewards. This is sufficient in traditional CB problems where the goal is to maximize total rewards, but may not be sufficient in other contexts, such as when CBs are used in adaptive treatment assignment with other goals or constraints [8, 25].

### v2

@@ 2 Related Work
Recently, researchers have attempted to integrate LLMs into social science studies, specifically utilizing LLMs to simulate human behaviours in social psychology experiments [17, 2] and predict their opinions [6, 40]. Additionally, a growing body of research has focused on evaluating LLM performance and biases through survey questions. Prior studies have examined how LLMs respond to multiple-choice questions from public opinion polls, revealing insights into their alignment with human responses and potential biases [39, 19, 18]. Santurkar et al. [39] focus on diverse demographic groups across the U.S., while Durmus et al. [19] focus on non-American demographics, both finding that LLMs’ responses to surveys show discrepancies with the collected answers from representative samples. Furthermore, Dominguez-Olmedo et al. [18] show that models’ responses are sensitive to ordering and labelling biases, and they do not contain the statistical signals typically found in human populations.
@@ 3 Preliminaries
In a multi-armed bandit problem, an agent, at each step $t\in[T]$, is presented with a set of $K$ possible arms and must choose an arm $k_{t}$. After the agent chooses, they receive a reward $r_{t,k_{t}}$ drawn from a reward distribution $R(k_{t})$, which is initially unknown. We focus on the stochastic setting, where each $r_{t,k}$ is sampled independently from $R(k)$. The goal of the agent is to maximize the total reward $\sum_{t=1}^{T}r_{t,k_{t}}$. To achieve this, the agent needs to balance the trade-off between exploration (i.e., trying different arms to gather more information about their reward distribution) and exploitation (i.e., choosing among the arms that appear to offer the highest reward based on the collected information so far).
Contextual Multi-Armed Bandits.
In contextual multi-armed bandits (CB), where at each time step $t\in[T]$, the agent receives side information $\phi_{t}\in\mathcal{C}$ which is called the context. The reward distribution of pulling arm $k\in[K]$ at time $t$ may depend on the context $\phi_{t}$, that is, $r_{t,k}\sim R(k,\phi_{t})$
@@
A policy $\pi$ is a map from histories $H_{t}=(\phi_{1},k_{1},r_{1,k_{1}},\ldots,\phi_{t-1},k_{t-1},r_{t-1,k_{t-1}},\phi_{t})$ of contexts, arms, and realized rewards to the next arm $\pi(H_{t})=k$ the agent will choose.
@@
The (cumulative) regret of an agent after $T$ steps choosing arms based on a policy $\pi$ relative to a policy $\pi^{\ast}$ is given by
$\operatorname{reg}_{\pi,\pi^{\ast}}=\mathbb{E}\left[\sum_{t=1}^{T}r_{t,k^{*}_{t}}\right]-\mathbb{E}\left[\sum_{t=1}^{T}r_{t,k_{t}}\right]$
where $k_{t}$ is chosen according to $\pi$ and $k^{\ast}_{t}$ is chosen according to $\pi^{\ast}$ and the expectations are taken with respect to any randomness in the rewards, contexts, and policy.
@@
In some of our applications, we will consider sleeping bandits, where at time $t$, the agent may choose between a subset of arms $\mathcal{A}_{t}\subseteq[K]$ [26]. In this case, the bandit algorithm needs to learn and determine which arms yield the best reward when they are available.
@@ 4.1 Problem Formulation
We consider a set of $n$ users, where the vector of features of each user $i\in[n]$ is denoted by $X_{i}$, which is sampled from a population $\mathcal{X}$. At each time step $t$, user $u_{t}$ is sampled from the set of $n$ users. The user’s context can be calculated using a mapping function $\phi:\mathcal{X}\rightarrow\mathcal{C}$ where $\mathcal{C}$ is the context space. There are $K$ different arms, each representing a potential recommendation for the user. For each user $u_{t}$ and arm $k$, there is a scalar reward $R(k,u_{t})$ following some unknown distribution. For brevity, we write $R_{t}$ henceforth. We note that there is an optimal policy $\pi_{o}$ which chooses the optimal arm $k^{\ast}_{t}=\operatorname*{arg\,max}_{k\in[K]}\mathbb{E}[R_{t}(k)]$ at each time step. We seek to optimize the following (expected, cumulative) regret:
$\pi^{*}=\arg\min_{\pi}\operatorname{reg}_{\pi,\pi_{o}}$
@@ 5 Experiments
We validate the effectiveness of our proposed approach through two sets of experiments. [^1] Code for both experiments is available at https://github.com/BorealisAI/jump-starting-bandits. In both sets of experiments, we adapt the LinUCB algorithm [13] to train our bandit. The first set of experiments is based on synthetic data, and the second set uses a real-world dataset derived from a choice-based conjoint analysis [28]. Experimental details are in Appendix C.
@@ 5.1 First Experiment: Personalized Email Campaign for Charity Donations
For this experiment, the ideal participants would be actual donors of our target charity. These users would be characterized by a rich set of attributes such as age, gender, location, occupation, hobbies, and their history of charitable donations. However, obtaining such comprehensive and detailed user data is not only challenging due to the sensitivity and privacy concerns associated with handling personal information but also financially and operationally demanding. Therefore, in this set of experiments, we propose an innovative alternative. Instead of relying on actual user datasets, we utilize a larger model’s responses as an oracle to pre-train our bandit on the responses of a smaller model. In this experiment, we rely on OpenAI GPT-4o model [1] as a surrogate to the real potential donors, and simulate their preferences.
@@ 5.2 Second Experiment: A Choice-Based Conjoint Analysis with Real-world Data
To train our model according to Algorithm 1, we generate 10,000 synthetic users, based on the set of features of the actual users. We use these users to pre-train our contextual bandit.
@@ 7 Limitations
Finally, our reward function estimation focuses on determining the rank order of rewards and not the magnitude of rewards. This is sufficient in traditional CB problems where the goal is to maximize total rewards, but may not be sufficient in other contexts, such as when CBs are used in adaptive treatment assignment with other goals or constraints [8, 25].